# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v1

@@ Abstract
footnotetext: ————————————————————footnotetext: * Equal contribution. random order.footnotetext: † Corresponding author.
@@ I Introduction
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, edge=-, opacity=1.0, line width=1.pt, , where level=1text width=12.0em,font=,edge=line width=1.pt, where level=2text width=12em,font=,edge=line width=1.pt, [Discrete Diffusion (M)LLMs, fill=Gray!10, edge=Black!60 [Mathematical Formulations (Sec. II), fill=lan!19, edge=lan!40 [Discrete Denoising Diffusion
Probabilistic Model (Sec. II-A), fill=lan!12, edge=lan!40 [Discrete Diffusion for Binary Variable [27],
@@
D3PM [13], leaf, text width=16em, fill=lan!12, edge=lan!40] ] [Reparameterized Discrete Diffusion
Model (Sec. II-B), fill=lan!12, edge=lan!40 [RDM [15], MLDM [16], MD4 [17], leaf, text width=16em, fill=lan!12, edge=lan!40] ] [Continuous Time Discrete Denoising
Models [28] (Sec. II-C), fill=lan!12, edge=lan!40 ] [Concrete Score (Sec. II-D), fill=lan!12, edge=lan!40 [CSM [29], DWDSE [16], Categorical Ratio
Matching Loss [30], RADD [22], leaf, text width=16em, fill=lan!12, edge=lan!40] ] [Discrete Flow Matching [24]
(Sec. II-E), fill=lan!12, edge=lan!40 ] [Block Diffusion Models [31]
(Sec. II-F), fill=lan!12, edge=lan!40 ] ] [Representative Models (Sec. III), fill=mul!19, edge=mul!40 [Discrete Diffusion Models around
1B (Sec. III-A), fill=mul!12, edge=mul!40 [ D3PM [13], DiffusionBERT [32], RDM [15],
@@
MD4 [17], UniDisc [38] , leaf, text width=16em, fill=mul!12, edge=mul!40] ] [Large Diffusion Language Models
(Sec. III-B), fill=mul!12, edge=mul!40 [ LLaDA [6], DIFFUSION-LLMs[39],
DiffuGPT & DiffuLLaMA[18], DREAM[7],
LLaDA 1.5 [40], TESS 2[41] , leaf, text width=16em, fill=mul!12, edge=mul!40] ] [Large Diffusion Multimodal Models
(Sec. III-C), fill=mul!12, edge=mul!40 [ Dimple [8], LaViDa [10], LLaDA-V [9] , leaf, text width=16em, fill=mul!12, edge=mul!40] ] [Large Unified Models (Sec. III-D), fill=mul!12, edge=mul!40 [ MMaDA [11], FUDOKI [42], Muddit [43] , leaf, text width=16em, fill=mul!12, edge=mul!40] ] ] [Training Techniques (Sec. IV), fill=uni!19, edge=uni!40 [Initialization Technique (Sec. IV-B), fill=uni!12, edge=uni!40 [ BERT Initialization [32], Autoregressive
Initialization[18, 7], Autoregressive-then
Diffusion Training [8] , leaf, text width=16em, fill=uni!12, edge=uni!40] ] [Complementary Masking Technique
[10] (Sec. IV-C), fill=uni!12, edge=uni!40 ] [Masking Scheduling Technique
(Sec. IV-D), fill=uni!12, edge=uni!40 [ Linear Schedule [13], Geometric Schedule [37],
Cosine Schedule [44], Token-wise Masking
Scheduling [32] , leaf, text width=16em, fill=uni!12, edge=uni!40] ] ] [Inference Techniques (Sec. V), fill=SkyBlue!19, edge=SkyBlue!40 [Unmasking Techniques (Sec. V-A), fill=SkyBlue!12, edge=SkyBlue!40 [ Metrics Used in Unmasking: Confidence,
Margin, Entropy [7], Confident Decoding [8],
@@
Continuous Time (Flow Matching)
Unmasking [24] , leaf, text width=16em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Remasking Techniques (Sec. V-B), fill=SkyBlue!12, edge=SkyBlue!40 [ Discrete Time Remasking [23], Continuous Time
(Flow Matching) Remasking [24] , leaf, text width=16em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Prefilling and Caching Technique
(Sec. V-C), fill=SkyBlue!12, edge=SkyBlue!40 [ Prefilling [8, 10], dKV-Cache [25],
dLLM-Cache [26] , leaf, text width=16em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Guidance Technique (Sec. V-D), fill=SkyBlue!12, edge=SkyBlue!40 [ Classifier-Free Guidance [19], Classifier
Guidance [45], Reward Guidance [41] , leaf, text width=16em, fill=SkyBlue!12, edge=SkyBlue!40] ] ] [Applications (Sec. VI), fill=YellowOrange!19, edge=YellowOrange!40 [Text Generation and Style Control
(Sec. VI-A), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffuSeq_StylePTB [46], REASONAUG [47],
SLD [48], DiffusPoll [49], PoetryDiffusion [50] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Text Editing and Summarization
(Sec. VI-B), fill=YellowOrange!12, edge=YellowOrange!40 [ EdiText [51], CrossMamba [52], DiffETM [53],
TermDiffuSum [54] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Sentiment Analysis and Data
Augmentation (Sec. VI-C), fill=YellowOrange!12, edge=YellowOrange!40 [ $\textrm{CDA}^{2}$ [55], DiffusionCLS [56], GDP [57] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Knowledge and Reasoning
(Sec. VI-D), fill=YellowOrange!12, edge=YellowOrange!40 [ DoT [58], DiffuCOMET [59], DPCL-Diff [60],
d1 [61] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Vision and Multimodal (Sec. VI-E), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffVLA [62], UDAN-CLIP [63], VPDD [64],
M2D2M [65], AR-Diffusion [66] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Biological and Drug Discovery
(Sec. VI-F), fill=YellowOrange!12, edge=YellowOrange!40 [ MolEditRL [67], CFP-Gen [68], TransDLM [69],
GenMol [70], DPLM-2 [71], PepTune [72] , leaf, text width=16em, fill=YellowOrange!12, edge=YellowOrange!40] ] ] ]
@@ II Mathematical Formulations
II-A Discrete Denoising Diffusion Probabilistic Model (D3PM)
@@
  • Uniform: $Q_{t}=(1-\beta_{t})I+\frac{\beta_{t}}{K}\mathbf{1}\mathbf{1}^{\top}$ yields a uniform stationary distribution.
  • Absorbing: Tokens either remain unchanged or are mapped to a special [MASK] token with probability $\beta_{t}$.
  • Discretized Gaussian: For ordinal data, transitions favor neighboring categories, imitating Gaussian kernels.
  • Embedding-based: Transition probabilities are based on token similarity in embedding space, guiding structured corruption.
@@
II-B Reparameterized Discrete Diffusion Model (RDM)
@@
$\log p_{\theta}(x_{0})\geq\mathcal{L}_{1}-\sum_{t=2}^{T}\mathcal{L}_{t}+C,$ (18)
@@
$\mathcal{L}_{t}$ $=\mathbb{E}_{x_{0},x_{1:T}}\Big{[}-\lambda_{t-1}^{(2)}\sum_{n=1}^{N}(1-b_{t,n})x_{0,n}^{\top}\log f(x_{t,n};\bm{\theta})\Big{]},$ (20)
where $\lambda_{t-1}^{(2)}=\frac{\alpha_{t-1}-\alpha_{t}}{1-\alpha_{t}}$. This loss corresponds to a reweighted cross-entropy loss evaluated only over noisy tokens. Such loss formulation is significantly simpler than the original variational bound and has become the standard training objective for subsequent large discrete diffusion models. Similar loss are also derived in [16, 17], with different simplification and reparameterization tricks.
II-C Continuous Time Discrete Denoising Models

### v2

@@ Abstract
footnotetext: ————————————————————footnotetext: * Equal contribution, random order.footnotetext: † Corresponding author.
@@ I Introduction
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, edge=-, opacity=1.0, line width=1.pt, , where level=1text width=12.0em,font=,edge=line width=1.pt, where level=2text width=12em,font=,edge=line width=1.pt, [Discrete Diffusion (M)LLMs, fill=Gray!10, edge=Black!60 [Mathematical Formulations (Sec. II), fill=lan!19, edge=lan!40 [Discrete Diffusion Model & Transi-
tion Matrix (Sec. II.A), fill=lan!12, edge=lan!40 [Discrete Diffusion for Binary Variable [27],
@@
D3PM [13], Roulette Diffusion [28], GIDD [29], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Reparameterized Discrete Diffusion
Model [15] (Sec. II.B), fill=lan!12, edge=lan!40 ] [Simplified Masked Diffusion
Model (Sec. II.C), fill=lan!12, edge=lan!40 [MLDM [16], MD4 [17], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Continuous Time Discrete Denoising
Models [30] (Sec. II.D), fill=lan!12, edge=lan!40 ] [Concrete Score (Sec. II.E), fill=lan!12, edge=lan!40 [CSM [31], DWDSE [16], Categorical Ratio Matching
Loss [32], RADD [22], TCSM [33], CEDD [28], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Discrete Flow Matching [24]
(Sec. II.F), fill=lan!12, edge=lan!40 ] [Block Diffusion Models [34]
(Sec. II.G), fill=lan!12, edge=lan!40 ] ] [Representative Models (Sec. III), fill=mul!19, edge=mul!40 [Discrete Diffusion Models around
1B (Sec. III.A), fill=mul!12, edge=mul!40 [ D3PM [13], DiffusionBERT [35], RDM [15],
@@
MD4 [17], UniDisc [41] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Diffusion Language Models
(Sec. III.B), fill=mul!12, edge=mul!40 [ LLaDA [6], DIFFUSION-LLMs[42], DiffuGPT &
DiffuLLaMA[18], DREAM[7], LLaDA 1.5 [43],
TESS 2[44] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Diffusion Multimodal Models
(Sec. III.C), fill=mul!12, edge=mul!40 [ Dimple [8], LaViDa [10], LLaDA-V [9] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Unified Models (Sec. III.D), fill=mul!12, edge=mul!40 [ MMaDA [11], FUDOKI [45], Muddit [46] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] ] [Training Techniques (Sec. IV), fill=uni!19, edge=uni!40 [Initialization Technique (Sec. IV.B), fill=uni!12, edge=uni!40 [ BERT Initialization [35], Autoregressive Initiali-
zation[18, 7], Autoregressive-then Diffusion
Training [8] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Complementary Masking Technique
[10] (Sec. IV.C), fill=uni!12, edge=uni!40 ] [Masking Scheduling Technique
(Sec. IV.D), fill=uni!12, edge=uni!40 [ Linear Schedule [13], Geometric Schedule [40], Cosine
Schedule [47], Token-wise Masking Scheduling [35] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Addressing Training-Testing
Discrepancy (Sec. IV.E), fill=uni!12, edge=uni!40 [ Two-Step Strategy [48] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Reweighting Technique (Sec. IV.F), fill=uni!12, edge=uni!40 [ MGDM [49] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Distillation (Sec. IV.G), fill=uni!12, edge=uni!40 [ Di4C [50] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] ] [Inference Techniques (Sec. V), fill=SkyBlue!19, edge=SkyBlue!40 [Unmasking Techniques (Sec. V.A), fill=SkyBlue!12, edge=SkyBlue!40 [ Metrics Used in Unmasking: Confidence, Margin,
Entropy [7], Confident Decoding [8],
@@
Continuous Time (Flow Matching) Unmasking [24] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Remasking Techniques (Sec. V.B), fill=SkyBlue!12, edge=SkyBlue!40 [ Discrete Time Remasking [23], Continuous Time
(Flow Matching) Remasking [24] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Prefilling and Caching Technique
(Sec. V.C), fill=SkyBlue!12, edge=SkyBlue!40 [ Prefilling [8, 10], dKV-Cache [25], dLLM-Cache [26],
DualCache [51] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Guidance Technique (Sec. V.D), fill=SkyBlue!12, edge=SkyBlue!40 [ Classifier-Free Guidance [19], Classifier Guidance [52],
Reward Guidance [44], EDLM [53] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] ] [Applications (Sec. VI), fill=YellowOrange!19, edge=YellowOrange!40 [Text Generation and Style Control
(Sec. VI.A), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffuSeq_StylePTB [54], DiffEmbed [55], SLD [56],
DiffusPoll [57], PoetryDiffusion [58] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Text Editing and Summarization
(Sec. VI.B), fill=YellowOrange!12, edge=YellowOrange!40 [ EdiText [59], CrossMamba [60], DiffETM [61],
TermDiffuSum [62] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Sentiment Analysis and Data
Augmentation (Sec. VI.C), fill=YellowOrange!12, edge=YellowOrange!40 [ $\textrm{CDA}^{2}$ [63], DiffusionCLS [64], GDP [65] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Knowledge and Reasoning
(Sec. VI.D), fill=YellowOrange!12, edge=YellowOrange!40 [ DoT [66], DiffuCOMET [67], DPCL-Diff [68],
d1 [69], MGDM [49] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Vision and Multimodal (Sec. VI.E), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffVLA [70], UDAN-CLIP [71], VPDD [72],
M2D2M [73], AR-Diffusion [74] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Biological and Drug Discovery
(Sec. VI.F), fill=YellowOrange!12, edge=YellowOrange!40 [ MolEditRL [75], CFP-Gen [76], TransDLM [77],
GenMol [78], DPLM-2 [79], PepTune [80] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] ] ]
@@ II Mathematical Formulations
II.A Discrete Diffusion Model and Transition Matrix
@@
  • Uniform: $Q_{t}=(1-\beta_{t})I+\frac{\beta_{t}}{K}\mathbf{1}\mathbf{1}^{\top}$ yields a uniform stationary distribution. The uniform transition matrix looks like
    $Q_{t}^{\text{uniform}}=\begin{bmatrix}1-\frac{K-1}{K}\beta_{t}&\frac{\beta_{t}}{K}&\cdots&\frac{\beta_{t}}{K}\\ \frac{\beta_{t}}{K}&1-\frac{K-1}{K}\beta_{t}&\cdots&\frac{\beta_{t}}{K}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\beta_{t}}{K}&\frac{\beta_{t}}{K}&\cdots&1-\frac{K-1}{K}\beta_{t}\end{bmatrix}.$ (7)
  • Absorbing: $Q_{t}=(1-\beta_{t})I+\beta_{t}\mathbf{1}e_{m}^{\top}$, where $e_{m}$ is a vector with a one on the absorbing state and zeros elsewhere. Tokens either remain unchanged or are mapped to a special [MASK] token with probability $\beta_{t}$. The absorbing transition matrix looks like
    $Q_{t}^{\text{absorb}}=\begin{bmatrix}1-\beta_{t}&0&\cdots&0&\beta_{t}\\ 0&1-\beta_{t}&\cdots&0&\beta_{t}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1-\beta_{t}&\beta_{t}\\ 0&0&\cdots&0&1\end{bmatrix}.$ (8)
  • Hybrid: Hybrid transition was initially discussed in [13], where multiple types of transition are combined to create more expressive diffusion processes. For example, the Routlette Diffusion in [28] and the Generalized Interpolating Discrete Diffusion (GIDD) [29] both study the linear combination of the absorbing transition and the uniform transition. A key motivation for such “Linear + Absorbing” transitions is to overcome a fundamental limitation of absorbing diffusion: once a token is denoised, it cannot be revised. Such “Linear + Absorbing” transitions address this by allowing previously denoised tokens to re-enter the diffusion process via the uniform component. This grants the model the ability to correct earlier generation errors during inference.
  • Discretized Gaussian:
    $[Q_{t}]_{ij}=\begin{cases}\frac{\exp\left(-\frac{4|i-j|^{2}}{(K-1)^{2}\beta_{t}}\right)}{\sum\limits_{n=-(K-1)}^{K-1}\exp\left(-\frac{4n^{2}}{(K-1)^{2}\beta_{t}}\right)}&\text{if }i\neq j\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&\text{if }i=j\end{cases}$
    This matrix simulates Gaussian diffusion, suitable for ordinal data. Each state $i$ is most likely to transition to nearby states $j$ with probabilities resembling a Gaussian kernel. Closer states receive higher probabilities, while distant states receive lower ones. Diagonal values are chosen to ensure that each row sums to 1, yielding a uniform stationary distribution.
  • Band-diagonal:
    $[Q_{t}]_{ij}=\begin{cases}\frac{1}{K\beta_{t}}&0<|i-j|\leq v\\ 0&|i-j|>v\\ 1-\sum\limits_{l\neq i}[Q_{t}]_{il}&i=j\end{cases}$
    Band-diagonal imposes local transitions only: state $i$ can only transition to its $v$-nearest neighbors on the ordinal scale. It biases the forward process toward small, local perturbations.
  • Embedding-based: Let $A$ be an adjacent matrix build from based on the similarity between tokens in the embedding space, in [13] the Embedding-based transition matrix is defined as
    $Q_{t}$ $=\exp(\alpha_{t}R)=\sum_{n=0}^{\infty}\frac{\alpha_{t}^{n}}{n!}R^{n},$ (9)
    $R_{ij}$ =$=\begin{cases}-\sum_{l\neq i}A_{il}&i=j\\ A_{ij}&i\neq j\end{cases}.$ (10)
    This transition matrix promotes a diffusion process where tokens are more likely to transition to others that are semantically or syntactically similar in the embedding space.
@@
II.B Reparameterized Discrete Diffusion Model
@@
$\log p_{\theta}(x_{0})\geq\mathcal{L}_{1}-\underbrace{\sum_{t=2}^{T}\mathcal{L}_{t}}_{\mathcal{L}_{T}}+C,$ (22)
@@
$\mathcal{L}_{t}$ $=\mathbb{E}_{x_{0},x_{1:T}}\Big{[}-\lambda_{t-1}^{(2)}\sum_{n=1}^{N}(1-b_{t,n})x_{0,n}\log[f_{\theta}(x_{t})]_{n}\Big{]},$ (24)
where $\lambda_{t-1}^{(2)}=\frac{\alpha_{t-1}-\alpha_{t}}{1-\alpha_{t}}$.
A central issue of RDM lies in its dependence on the ground truth $x_{0}$ to compute the backward transition probabilities Eq. 17. However, in the inference stage, $x_{0}$ is unknown, making it infeasible to directly evaluate the indicator function $b_{t}$ required for determining the appropriate transition path. To overcome this limitation, the authors propose a recursive approximation for computing $b_{t}$ by utilizing the Bernoulli routing variables $v$. Beginning with $b_{T}=0$, which assumes the initial sequence is fully noisy, the clean token set is recursively updated via:
$b_{t-1,n}=\left(b_{t,n}\land v_{t-1,n}^{(1)}\right)\lor v_{t-1,n}^{(2)},$ (25)
where $\land$ and $\lor$ denote logical conjunction and disjunction, respectively.
II.C Simplified Masked Diffusion Model
A widely adopted class of discrete diffusion models is based on an absorbing state and is often referred to as the Masked Diffusion Model. Both [17, 16] introduced simplifications to the diffusion process and the corresponding training objective for masked diffusion, yielding improved performance and computational efficiency.
In masked diffusion, the forward process progressively replaces input tokens with a special [MASK] token. Once a token is masked, it remains in that state throughout the remainder of the process, making the [MASK] token an absorbing state. At each time step $t$, the forward transition for a token $x$ is defined as:
$q(x_{t}\mid x_{0})=\text{Cat}(x_{t};\alpha_{t}x_{0}+(1-\alpha_{t})m),$ (26)
where $m$ is the one-hot vector corresponding to the [MASK] token, and $\alpha_{t}\in[0,1]$ is the monotonically decreasing schedule such that $\alpha_{0}\approx 1$ and $\alpha_{T}=0$.
The reverse process aims to denoise the masked sequence by substituting [MASK] tokens with predicted tokens. Importantly, unmasked tokens are carried out unchanged throughout the denoising process. The reverse posterior at a previous time $s$ conditioned on $x_{t}$ and $x_{0}$ is given by:
$q(x_{s}\mid x_{t},x_{0})=$
$\begin{cases}\text{Cat}(x_{s};x_{t}),&\text{if }x_{t}\neq m,\\ \text{Cat}\left(x_{s};\dfrac{(1-\alpha_{s})m+(\alpha_{s}-\alpha_{t})x_{0}}{1-\alpha_{t}}\right),&\text{if }x_{t}=m.\end{cases}$ (27)
This formulation reflects two key properties of the masking process: (1) If the current token $x_{t}$ is not masked, the posterior is deterministic: $x_{s}=x_{t}$. (2) If the token is masked, then the posterior is a linear interpolation between the mask vector $m$ and the clean token $x_{0}$, scaled by the noise schedule parameters $\alpha_{s}$ and $\alpha_{t}$.
Let $f_{\theta}(x_{t})$ be the neural network output predicting the original token $x_{0}$ from the noisy input $x_{t}$. The above reverse transition is rewritten as:
$p_{\theta}(x_{s}\mid x_{t})=\begin{cases}\text{Cat}(x_{s};x_{t}),&\text{if }x_{t}\neq m,\\ \text{Cat}\left(x_{s};\frac{(1-\alpha_{s})m+(\alpha_{s}-\alpha_{t})f_{\theta}(x_{t})}{1-\alpha_{t}}\right),&\text{if }x_{t}=m.\end{cases}$ (28)
The variational lower bound for discrete diffusion can be simplified using the above formulation, leading to an final loss in the form of
$\mathcal{L}$ $=\sum_{t=2}^{T}\mathbb{E}_{x_{0},x_{1:T}}\Big{[}-\frac{\alpha_{t-1}-\alpha_{t}}{1-\alpha_{t}}\sum_{n=1}^{N}\delta_{m}(x_{t,n})x_{0,n}\log[f_{\theta}(x_{t})]_{n}\Big{]},$ (29)
where $\delta_{m}(x_{t,n})$ denotes the indicator function, $\delta_{m}(x_{t,n})=1$, if the n-th token of $x_{t}$ is a masked token, otherwise, $\delta_{m}(x_{t,n})=0$. This loss is similar to Eq. 24 in the form, but the masked diffusion model eliminates the need to rely on ground truth for computing the indicator function; it can instead be determined based on the mask tokens.
By defining the discrete time series as $10,\frac{1}{T},\ldots,1-\frac{1}{T},$ 1 and letting $T\to\infty$, both works [16, 17] extend the above loss Eq. 29 to the continuous-time setting:
$\mathcal{L}$ $=\int_{0}^{1}\mathbb{E}_{x_{0:1}}\Big{[}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{n=1}^{N}\delta_{m}(x_{t,n})x_{0,n}\log[f_{\theta}(x_{t})]_{n}\Big{]}dt.$ (30)
This loss corresponds to a reweighted cross-entropy loss evaluated only over masked tokens. Such loss formulation is significantly simpler than the original variational bound and has become the standard training objective for subsequent large discrete diffusion models.
II.D Continuous Time Discrete Denoising Models